# Semi-supervised and unsupervised learning for health indicator extraction from guided waves in aerospace composite structures

### v1

@@ 2.6.2 DTC-VAE
The DTC-VAE model is optimised by minimising the loss function, $\mathcal{L}_{DTC-VAE}$, which is a summation of the Kullback-Leibler divergence ($\mathcal{L}_{KL}$), reconstruction loss ($\mathcal{L}_{RE}$) and monotonicity constraint loss ($\mathcal{L}_{Mo}$):
@@
where $\alpha$, $\beta$, and $\gamma$ are their respective hyperparameters.
Firstly, $\mathcal{L}_{KL}$ quantifies the difference between the prior distribution and the posterior distribution approximated by the encoder, $\phi(z_{j}|x_{j})$. In this case, the prior is assumed to be a simple standard normal distribution $p(z)\sim N(0,1)$, simplifying $\mathcal{L}_{KL}$ to the form presented in Equation 18:
$\mathcal{L}_{KL}(\phi,\psi)=-\frac{1}{2}(1+\log(\sigma^{2})-\sigma^{2}-\mu^{2})$ (18)
which only includes the standard deviation $\sigma$ and mean $\mu$ of the encoder distribution.
@@
Finally, the latent variable is prompted to follow a monotonic trend by adding the monotonicity constraint loss, $\mathcal{L}_{Mo}$. This was proposed by Qin et al. [55], ensuring the extracted HIs followed a monotonic trend by adding a degradation-trend constraint, hence the name DTC-VAE. While the other loss terms are common for VAE architectures, this one is unique to the DTC-VAE model and has a direct effect on the fitness of the output, which is expressed as follows:
@@
where $z_{j}^{m}$ represents the latent variable for specimen $m$ at timestep $j$, while the rate of degradation $r$ is a hyperparameter ensuring that $z_{j}^{m}>\left(z_{j-1}^{m}+r\right)$, which controls the magnitude of change of a HI between timesteps. Its value was determined from the original DTC-VAE study [55], in essence, a randomly selected constant in the range $(9,10)$.
@@
The original HIs criteria functions could not be adapted for gradient descent, given the fact that they are not differentiable and require comparison of latent outputs across specimens, inaccessible during model training. In the case of monotonicity, originally presented in Equation 1, an alternate formulation $\mathcal{L}_{Mo}$ has been implemented, which ensures high monotonicity scores and is compatible with the training process. Ideally, all prognostic criteria (Mo, Pr, Tr) shall be implemented into the loss function in future to maximise fitness.
@@
A common issue in VAEs is posterior collapse, where the latent variables fail to encode meaningful information because the decoder learns to reconstruct inputs directly and the posterior distribution collapses to the prior. This typically occurs when the KL divergence term dominates the loss [64]. In this work, the risk of collapse is mitigated through optimisation of the loss weights $\alpha$, $\beta$, and $\gamma$, following Qin et al. [55], and by the inclusion of the monotonicity constraint $\mathcal{L}_{Mo}$, which forces the latent variables to capture degradation trends instead of defaulting to the prior.
@@
DTC-VAE architecture and hyperparameter optimisation: The DTC-VAE model consists of 3 components: the encoder, latent space and decoder. The encoder’s only hidden layer uses a sigmoid activation function and has a varying number of neurons, kept as a hyperparameter. This is followed by two output layers of size 1, representing the mean and log variance of the latent space. As explained by Equation 16, the reparameterization trick is used to allow for gradient calculation and thus backpropagation. The decoder acts as a mirror to the encoder, it takes the latent variable as input to a hidden layer with sigmoid activation functions. This hidden layer has the same number of neurons as in the encoder. It is followed by an output layer the same size as the input with linear activation function, which returns the reconstructed input. All weights are initialised with uniform Glorot (Xavier) initialisation [65] due to its state-of-the-art performance with sigmoid activation functions and improved stability. Adam is used as the optimiser.
The hyperparameters were optimised over the space detailed in Table 3, using the Bayesian optimisation algorithm maximizing the $F_{all}$ score as for Diversity-DeepSAD. From the preprocessed data and the optimised hyperparameters, a model was trained for every fold and frequency.
@@ 2.6.3 Ensemble learning model
The weighted averaging ensemble (WAE) model was implemented to fuse the HI results of the different excitation GW frequencies. This is done to improve performance by reducing instability, and removing the need to select and rely on only one frequency. It was chosen to use the normalised fitness scores of the averaged HIs as weights $\tilde{\omega}_{f}$ for the same frequency, as this method was most successfully implemented previously [15]. This method is described by Equation 21, with the constants $\omega_{f}$ adjusted depending on their fitness, producing a new weighted-average HI value $\bar{y}$.
$\bar{y}=\sum_{f=1}^{F}\tilde{\omega}_{f}\bar{y}^{f}$ (21a)
@@ 3 Results
In this section, signal processing methods are first compared based on the HIs criteria calculated from the statistical features extracted for each method. Following this, the results of the semi-supervised Diversity-DeepSAD model and the unsupervised DTC-VAE model are presented and compared using the HIs criteria across all specimens $F_{all}$ (training, validation, and test) as well as for test specimens alone $F_{test}$. Beyond comparing the two deep learning models, the analysis also examines the impact of different excitation GW frequencies, the fusion of all excitation GW frequencies, and the selection of signal processing methods (focusing on the two top-performing ones).
@@ 3.1 Signal Processing
Figure 6 shows the $F_{all}$ scores averaged across all paths and frequencies for each of the extracted statistical features. It is clear that all signal processing methods contain some features that perform poorly when evaluated for fitness. In order to discard the low-performing features, the set is reduced to only those performing above average ($\mu=1.614$), which is indicated by the red horizontal line in Figure 6. The mean score of features extracted from each SP method is presented in Table 4. It can be seen that after the feature reduction, FFT and HT yield the highest-performing features, which will be used for further individual use in the following AI models.
@@ 3.2 Diversity-DeepSAD
The Diversity DeepSAD model was trained across all combinations of frequencies and data folds (i.e., training and test divisions), with WAE implemented as described. The HIs constructed by Diversity DeepSAD upon FFT and HT features are shown in Figure 7 and Figure 8, respectively. For better comparison, the HIs are displayed over the normalised lifetime of specimens, ranging from 0% to 100%. The resulting fitness scores of the HIs are reported in Table 5 and Table 6. These tables include $F_{all}$ scores, from Equation 6, and $F_{test}$ scores, from Equation 7, for HIs generated from both FFT and HT features.
@@
Diversity-DeepSAD results show high fitness scores for both FFT and HT, with mean $F_{test}$ scores of 2.27 and 2.13 respectively, with $F_{all}$ scores averaging slightly higher at 2.35 and 2.22. This implies a general higher performance of features from the FFT than those from the HT. No fold performed consistently worst than the others, although average results from test Sample 1 were consistently highest. Inspecting the graph, this may be due to the invariably conforming behaviour of this sample, aiding in high $F_{test}$ scores. All frequencies performed similarly, with $50\text{\,}\mathrm{kHz}$ producing marginally better average $F_{all}$ results for both signal processing methods.
@@ 3.4 Weighted Average Ensemble
The stability of fused results is also greatly improved, with deviations between folds smaller than within individual frequencies. For both models, but still most noticeably DTC-VAE, the fused graphs are visually more consistent between folds than unfused. Between seeds, the standard deviations of fitness for the fusion are always below 0.1, with one exception in Diversity-DeepSAD fold 4. The average standard deviation for Diversity-DeepSAD-Fusion scores is half that of unfused, 0.05 compared to 0.10 for $F_{all}$ and 0.06 to 0.12 for $F_{test}$. For DTC-VAE, the reduction is better than half at 0.04 from 0.09 and 0.05 from 0.11 respectively. These low standard deviations indicate that HIs fused using WAE of multiple guided wave excitation frequencies are more robust. This is because different GW excitation frequencies can capture various damage modes in structures, and the ensemble approach reduces susceptibility to errors in individual constituent models.
@@ 3.5 Model comparison
The results reflect the fact that trendability was the lowest performing prognostic criterion for both models, as can be seen by the inconsistency of shapes between specimens in the HI graphs. This could be due to significant uncertainties introduced during the manufacturing and fatigue experiments, as explained in Subsection 2.1. It shows that the developed framework satisfies the first two fact-based criteria while striving to fulfill the aspirational trendability metric to capture the overall trend. Both models achieved moderate scores, with Diversity-DeepSAD scoring an average $Tr$ score of 0.59 and 0.50, and DTC-VAE scoring 0.68 and 0.52, for FFT and HT features respectively.

### v2

@@ 3.6.2 DTC-VAE
The DTC-VAE model is optimised by minimising the loss function, $\mathcal{L}_{DTC-VAE}$, which is a summation of the Kullback-Leibler divergence ($\mathcal{L}_{KL}$) [kingma2014autoencoding], reconstruction loss ($\mathcal{L}_{RE}$) and monotonicity constraint loss ($\mathcal{L}_{Mo}$):
@@
where $\alpha$, $\beta$, and $\gamma$ are their respective hyperparameter weights.
@@
Finally, the latent variable is prompted to follow a monotonic trend by adding the monotonicity constraint loss, $\mathcal{L}_{Mo}$. This was proposed by Qin et al. [Qin2022], ensuring the extracted HIs followed a monotonic trend by adding a degradation-trend constraint, hence the name DTC-VAE. While the other loss terms are common for VAE architectures, this one is unique to the DTC-VAE model and has a direct effect on the fitness of the output. The monotonicity constraint loss is expressed as follows:
@@
where $z_{j}^{m}$ represents the latent variable for specimen $m$ at timestep $j$, while the rate of degradation $r$ is a hyperparameter ensuring that $z_{j}^{m}>\left(z_{j-1}^{m}+r\right)$, which controls the magnitude of change of an HI between timesteps. Its value was determined from the original DTC-VAE study [Qin2022], in essence, a randomly selected constant in the range $(9,10)$.
@@
The original HIs criteria functions could not be adapted for gradient descent, given the fact that they are not differentiable and require comparison of latent outputs across specimens, inaccessible during model training. In the case of monotonicity, originally presented in Equation 1, an alternate formulation $\mathcal{L}_{Mo}$ has been implemented, which ensures high monotonicity scores and is compatible with the training process. Ideally, all prognostic criteria ($Mo$, PrPr, TrTr) shall be implemented into the loss function in future to maximise fitness.
@@
A common issue in VAEs is posterior collapse, where the latent variables fail to encode meaningful information because the decoder learns to reconstruct inputs directly and the posterior distribution collapses to the prior. This typically occurs when the KL divergence term dominates the loss [Lucas2019]. In this work, the risk of collapse is mitigated through optimisation of the loss weights $\alpha$, $\beta$, and $\gamma$, following Qin et al. [Qin2022], and by the inclusion of the monotonicity constraint $\mathcal{L}_{Mo}$, which forces the latent variables to capture degradation trends instead of defaulting to the prior.
@@
DTC-VAE architecture and hyperparameter optimisation: The DTC-VAE model consists of 3 components: the encoder, latent space and decoder. The encoder’s only hidden layer uses a sigmoid activation function and has a varying number of neurons, kept as a hyperparameter. This is followed by two output layers of size 1, representing the mean and log variance of the latent space. As explained by Equation 15, the reparameterisation trick is used to allow for gradient calculation and thus backpropagation. The decoder acts as a mirror to the encoder, it takes the latent variable as input to a hidden layer with sigmoid activation functions. This hidden layer has the same number of neurons as in the encoder. It is followed by an output layer the same size as the input with linear activation function, which returns the reconstructed input. All weights are initialised with uniform Glorot (Xavier) initialisation [Xavier] due to its state-of-the-art performance with sigmoid activation functions and improved stability. Adam is used as the optimiser.
The hyperparameters were optimised over the space detailed in Table 3, again using Bayesian optimisation with Gaussian processes and $F_{all}$ as the objective. For each frequency and feature type (SP method), the optimiser was run for 40 evaluations: the first 10 corresponded to random samples from the hyperparameter space, and the remaining 30 were guided by the Gaussian-process surrogate. At each evaluation, DTC-VAE was trained on the four training specimens of the current fold, and $F_{all}$ was computed from the resulting HIs of these four specimens only. The held-out panel did not contribute to the optimisation objective.
Convergence behaviour was similar to that of Diversity-DeepSAD: the best observed $F_{all}$ typically stabilised after approximately 15 iterations, with subsequent iterations exploring hyperparameters within a narrow band of performance. The final configuration (per frequency and method) was selected as the hyperparameter set achieving the highest $F_{all}$ across all iterations and retrained on the full training set before computing the test metrics. The sensitivity analysis in Section 3.7, Section Subsection 4.6 and B confirms that the selected loss-weight combinations lie inside broad plateaus of high performance, rather than isolated optima.
@@ 3.6.3 Ensemble learning model
The weighted averaging ensemble (WAE) model was implemented to fuse the HI results of the $N_{f}$ different excitation GW frequencies. This is done to improve performance by reducing instability, and removing the need to select and rely on only one frequency. It was chosen to use the normalised fitness scores of the averaged HIs as weights $\tilde{\omega}_{f}$ for the same frequency, as this method was most successfully implemented previously [Moradi2024]. This method is described by Equation 19, with the constants $\omega_{f}$ adjusted depending on their fitness, producing a new weighted-average HI value $\bar{y}$.
$\bar{y}=\sum_{f=1}^{N_{f}}\tilde{\omega}_{f}\bar{y}^{f}$ (19a)
@@
To summarise the framework, Algorithm 1 presents the full HI generation workflow. It links SP and feature selection with per-frequency model training under leave-one-unit-out cross-validation, prognostic metric computation across random initialisation seeds, and the final weighted frequency-fusion stage. The information pipeline during this process is also shown visually in Figure 7. Multi-frequency GW measurements are first processed via feature extraction methods, including the HT and FFT, to obtain informative signal representations. These features are then used as input to the proposed models, augmented Diversity-DeepSAD and DTC-VAE, to generate per-frequency HIs. Frequency fusion, performed using WAE, integrates the per-frequency HIs into a unified HI for each model-transform combination.
Figure 7: Schematic of the frequency-fusion pipeline, where multi-frequency guided waves are transformed into FFT- and HT-based features to generate per-frequency HIs, which are fused using prognostic-criteria–optimised weights.
Input: Guided wave measurements for all specimens and excitation frequencies
Output: Per-frequency and fused HI trajectories with associated prognostic metrics
1exStage 2-3: Feature extraction and selection
foreach GW measurement (specimen, frequency, path, timestep) do
  apply SP methods (FFT, STFT, EMD, HT)
  extract statistical features
foreach candidate feature do
 aggregate per timestep and specimen; compute MoMo, PrPr, TrTr and $F_{all}$ for feature
retain features with fitness above the mean
select the best-performing SP methods (i.e. FFT, HT)
1exStage 4a: Per-frequency HI generation
foreach selected SP method do
 foreach GW excitation frequency $f$ do
    foreach LOOCV fold (choose one specimen as test, remaining specimens as train) do
        build training and test feature sets for all timesteps
        foreach model $\in$ {Diversity-DeepSAD, DTC-VAE} do
           fine-tune hyperparameters by Bayesian optimisation, maximising $F_{all}$ given only training specimens
           foreach random seed number do
              train the fine-tuned model using the random seed number and generate per-frequency HIs $y^{m,f}_{(\text{seed})}(t)$ for all specimens
              compute MoMo, PrPr, TrTr, $F_{all}$ and $Mo_{test}$, $Pr_{test}$, $F_{test}$ for this combination (model, SP, $f$, fold, seed) and store the non-fused HIs and metrics
          aggregate prognostic metrics across seeds (mean and standard deviation) for this combination (model, SP, $f$, fold) for reporting
1exStage 4b: Frequency fusion (WAE)
foreach selected SP method do
 foreach model $\in$ {Diversity-DeepSAD, DTC-VAE} do
    foreach LOOCV fold do
       foreach random seed number do
           apply WAE to per-frequency HIs $y^{m,f}_{(\text{seed})}(t)$ across all GW excitation frequencies $f$ to obtain fused HIs $\bar{y}^{m}_{(\text{seed})}(t)$ for all specimens
           compute MoMo, PrPr, TrTr, $F_{all}$ and $Mo_{test}$, $Pr_{test}$, $F_{test}$ for this fused combination (model, SP, fold, seed) and store fused HIs and metrics
       aggregate fused prognostic metrics across seeds (mean and standard deviation) for this combination (model, SP, fold)
return Per-frequency and fused HIs together with their prognostic metrics
Algorithm 1 Overview of HI generation and frequency-fusion framework
3.7 Hyperparameter sensitivity analysis
In addition to Bayesian optimisation, a sensitivity analysis was performed to assess the robustness of Diversity-DeepSAD and DTC-VAE to their hyperparameters, and quantify how perturbations around the nominal hyperparameters affect the resulting HIs.
For Diversity-DeepSAD, the analysis focused on the three weighting hyperparameters in Equation 11 and Equation 12: the $L_{2}$ regularisation weight $\nu$, the auxiliary label weighting $\eta$, and the diversity term weight $\lambda$. Other hyperparameters were fixed to their per-frequency optimal values as baseline. A three-dimensional grid search was then performed over $\nu$, $\eta$ and $\lambda$ on a logarithmic scale, $\{10^{-3},10^{-2},10^{-1},1,10\}$ for each parameter. For every triplet $(\nu,\eta,\lambda)$ and each excitation frequency, the model was retrained using the same cross-validation folds as in the main experiments, and the HIs were recomputed for all specimens. The prognostic criteria in Section 3.3 were then evaluated to obtain both $F_{all}$ and $F_{test}$, and the scores were aggregated by averaging across folds and excitation frequencies. This allowed response surfaces $F_{all}(\eta,\nu|\lambda)$ and $F_{test}(\eta,\nu|\lambda)$ to be constructed for each value of $\lambda$, allowing the identification of regions of stable high performance and the verification of the configuration used in the main study.
As for DTC-VAE, a similar procedure was followed for the loss weights $\alpha$, $\beta$ and $\gamma$ in Equation 16. Similarly, other hyperparameters were fixed to optimal parameters found during Bayesian optimisation to create a per-frequency baseline, allowing a comparison where only the loss weights were altered. A regular grid was then sampled over $(\alpha,\beta,\gamma)$ beyond the ranges reported in Table 3. Again, response surfaces $F_{all}(\alpha,\beta|\gamma)$ and $F_{test}(\alpha,\beta|\gamma)$ were plotted for each $\gamma$, demonstrating the effect of each hyperparameter and justifying the convergence and stability in the given hyperparameter space.
@@ 4 Results
In this section, SP methods are first compared based on the HIs criteria calculated from the statistical features extracted for each method. Following this, the results of the semi-supervised Diversity-DeepSAD model and the unsupervised DTC-VAE model are presented and compared using the HIs criteria across all specimens $F_{all}$ (training, validation, and test) as well as for test specimens alone $F_{test}$. Beyond comparing the two deep learning models, the analysis also examines the impact of different excitation GW frequencies, the fusion of all excitation GW frequencies, and the selection of SP methods (focusing on the two top-performing ones).
@@ 4.1 Signal processing
Figure 8 shows the $F_{all}$ scores averaged across all paths and frequencies for each of the extracted statistical features. It is clear that all SP methods contain some features that perform poorly when evaluated for fitness. In order to discard the low-performing features, the set is reduced to only those performing above average ($\mu=1.614$), which is indicated by the red horizontal line in Figure 8. The mean score of features extracted from each SP method is presented in Table 4. It can be seen that after the feature reduction, FFT and HT yield the highest-performing features, which are used in the subsequent AI models.
@@ 4.2 Diversity-DeepSAD
The Diversity-DeepSAD model was trained across all combinations of frequencies and data folds (i.e. training and test divisions), with WAE implemented as described. The HIs constructed by Diversity-DeepSAD upon FFT and HT features are shown in Figure 9 and Figure 10, respectively. For better comparison, the HIs are displayed over the normalised lifetime of specimens, ranging from 0% to 100%. The resulting fitness scores of the HIs are reported in Table 5 and Table 6. These tables include $F_{all}$ scores, from Equation 6, and $F_{test}$ scores, from Equation 7, for HIs generated from both FFT and HT features.
Diversity-DeepSAD results show high fitness scores for both FFT and HT, with mean $F_{test}$ scores of 2.27 and 2.13 respectively, with $F_{all}$ scores averaging slightly higher at 2.35 and 2.22. This implies a general higher performance of features from the FFT than those from the HT. No fold performed consistently worse than the others, although average results from test Sample 1 were consistently highest. Inspecting the graph, this may be due to the invariably conforming behaviour of this sample, aiding in high $F_{test}$ scores. All frequencies performed similarly, with $50\text{\,}\mathrm{kHz}$ producing marginally better average $F_{all}$ results for both SP methods.
@@ 4.4 Weighted Average Ensemble
The stability of fused results is also greatly improved, with deviations between folds smaller than within individual frequencies. For both models, but still most noticeably DTC-VAE, the fused graphs are visually more consistent between folds than unfused. Between seeds, the standard deviations of fitness for the fusion are always below 0.1, with one exception in Diversity-DeepSAD fold 4. The average standard deviation for Diversity-DeepSAD-Fusion scores is half that of unfused, 0.05 compared to 0.10 for $F_{all}$ and 0.06 to 0.12 for $F_{test}$. For DTC-VAE, the reduction is better than half at 0.04 from 0.09 and 0.05 from 0.11 respectively. These low standard deviations indicate that HIs fused using WAE of multiple GW excitation frequencies are more robust. This is because different GW excitation frequencies can capture various damage modes in structures, and the ensemble approach reduces susceptibility to errors in individual constituent models.
@@ 4.5 Model comparison
The results reflect the fact that trendability was the lowest performing prognostic criterion for both models, as can be seen by the inconsistency of shapes between specimens in the HI graphs. Both models achieved moderate scores, with Diversity-DeepSAD scoring an average $Tr$ score of 0.59 and 0.50, and DTC-VAE scoring 0.68 and 0.52, for FFT and HT features respectively. As explained in Subsection 3.1, each specimen had differences in manufacturing and the presence or absence of manufacturing disbonds or impact events. These factors significantly impact GW propoagation characteristics, and it is therefore physically justified that the $Tr$ score produced was lower than that for $Mo$ or $Pr$.
@@
4.6 Sensitivity analysis of hyperparameters
The hyperparameter sensitivity analysis confirmed that the conclusions drawn in the previous subsections are not the result of fragile or excessively fine-tuned configurations. For both Diversity-DeepSAD and DTC-VAE, the response surfaces constructed for $F_{all}$ and $F_{test}$ show broad regions of consistently high performance around the nominal hyperparameters obtained via Bayesian optimisation.
(a) Diversity-DeepSAD using FFT features.
(b) Diversity-DeepSAD using HT features.
Figure 13: Hyperparameter sensitivity analysis of augmented Diversity-DeepSAD over $(\nu,\eta,\lambda)$, for the two SP methods.
Regarding Diversity-DeepSAD, the surfaces $F_{all}(\eta,\nu|\lambda)$ and $F_{test}(\eta,\nu|\lambda)$ for $\lambda$ of 0.001 and 0.01 in Figure 13, with full plots in B, revealed that performance for both features is governed primarily by the diversity weight $\lambda$, with some sensitivity to the auxiliary label weight $\eta$ and weaker dependence on the regularisation weight $\nu$. For small values of $\lambda$, a plateau of high fitness values was observed, with $F_{all}$ and $F_{test}$ remaining close to their optima. Only very large $\nu$ values systematically degraded performance, showing that the regularisation term can have a significant effect. The large dependence on $\lambda$ confirmed that the diversity term is essential for producing smooth, monotonic HIs, while the regularisation refined performance. Overall, the configuration used ($\eta=10,\ \nu=10,\ \lambda=0.001$) lies well within the stable high-performance region and is therefore fully justified.
Regarding DTC-VAE, the response surfaces $F_{all}(\alpha,\beta|\gamma)$ and $F_{test}(\alpha,\beta|\gamma)$ exhibited a stronger stability across all hyperparameters, with two examples shown in Figure 14 and all figures being included in B. Increases beyond the optimised regions did lead to gradual changes in fitness. Very small $\gamma$ values saw the greatest decrease in fitness, highlighting the importance of the monotonicity constraint in constructing reliable HIs. Furthermore, for small $\gamma$, increases in $\beta$ beyond the optimised region led to reduced fitness, consistent with an overly strong reconstruction loss term dominating the objective, reducing the effective influence of the KL term and thus degrading the learned latent structure. However, this was not observed when the monotonicity constraint was dominant, showing improved stability. In all cases, the ranges optimised in the main experiments fell inside a stable high-fitness region, rather than at a sharp optimum.
(a) DTC-VAE using FFT features.
(b) DTC-VAE using HT features.
Figure 14: Hyperparameter sensitivity analysis of DTC-VAE over $(\alpha,\beta,\gamma)$, for the two SP methods.
Overall, the sensitivity analysis shows that both models maintain high $F_{all}$ and $F_{test}$ scores across a non-trivial neighbourhood of their nominal hyperparameters. This demonstrates that the comparative advantages of FFT over HT, and of DTC-VAE over Diversity-DeepSAD, arise from genuinely robust model behaviour rather than from narrowly tuned hyperparameter settings.